OpenStudy (anonymous):

Find the slope of the tangent to the curve y = 3x^4 - 4x at x = 4 And Find the equation of tangent to the curve given by x = a sin^3 t, y = bcos^3 t at a point where t = pi/2

zepdrix (zepdrix):

Hmm I was trying to figure out that second one... So like....The slope of the tangent line is given by \(\Large\rm \dfrac{dy}{dx}\). We have parametric equations so for our dy/dx we have,\[\Large\rm \frac{dy}{dx}=\frac{(dx/dt)}{(dy/dt)}\]yes?

zepdrix (zepdrix):

Woops I think I wrote that upside down. :P

zepdrix (zepdrix):

\[\Large\rm \frac{dy}{dx}=\frac{(dy/dt)}{(dx/dt)}\]Ya ya ya, my bad.

zepdrix (zepdrix):

Take your derivatives, simplify it all down and you should get something like this, \[\Large\rm \frac{dy}{dt}=-\frac{b}{a}\cot(t)\]Then evaluate it at \(\Large\rm t=\frac{\pi}{2}\) to get the `slope` of your tangent line. We're trying to form a line: \(\Large\rm y=mx+b\).

zepdrix (zepdrix):

I think you can handle those derivatives in the middle, yes? c: Those shouldn't give you too much trouble. Just a little chain rule going on. How bout the first problem you posted? Stuck on that one also?

OpenStudy (sweetnikkie072):

If u messed up with derivatives..(as i am too)..i know how to do it algebraically.. when x= 4 for tangent slope...(instantaneous) you plug in the very next or before number like if it asked for x=4 u can plug in either x=3.9999 or x=4.0001 if it has asked for x=1(for example) then for the second value u can either plug in x=0.999 or x=1.001 now when you have the values of y1 and y2 use this formula slope (m)=(y2 - y1)/(x2-x1) that 'l give u your answer! hope this helps!